Raymond Supply, a national hardware chain, is considering purchasing a smaller chain, Strauss & Glazer Parts (SGP). Raymond's analysts project that the merger will result in the following incremental free cash flows, tax shields, and horizon values (in millions):

Year

1

2

3

4

FCF

$2

$3

$4

$7

Unlevered Horizon value

$75

Tax shield

$1

$1

$2

$3

Horizon value of tax shield

$32

Assume that all cash flows occur at the end of the year. SGP is currently financed with 30% debt at a rate of 8%. The acquisition would be made immediately, and if it is undertaken, SGP would retain its current $20 million of debt and issue enough new debt to continue at the 30% target level. The interest rate would remain the same. SGP's pre-merger beta is 1.5, and its post-merger tax rate would be 40%. The risk-free rate is 5% and the market risk premium (=ErM - rRF) is 6%.

What is the unlevered cost of equity of SGP?

What is the unlevered value of operations of SGP?

What is the value of tax shield of SGP?
Answer #1

All financials below are in $ mn

What is the unlevered cost of equity of SGP?

Levered beta, BL = 1.5,

Levered cost of equity, Kel = Risk free rate + BL x market risk premium = 5% + 1.5 x 6% = 14.00%

Proportion of debt, Wd = D / (D + E) = 30%; Proportion of equity, We = 1 - Wd = 1- 30% = 70%

Cost of debt, Kd = 8%

Hence, unlevered cost of equity = We x Kel + Wd x Kd = 70% x 14% + 30% x 8% = 12.20%

-------------------------------

What is the unlevered value of operations of SGP?

Please see the working below. Please pay attention to the second column, it explains how each row has been caalculated.

Year N 1 2 3 4
FCF C               2.00               3.00               4.00               7.00
Unlevered Horizon value HV             75.00
Total Cash flows CF = C + HV               2.00               3.00               4.00             82.00
Discount factor (Unlevered cost of equity) Keu (calculated in first part) 12.20%
Discount factor DF = (1 + Keu)-N           0.8913           0.7944           0.7080           0.6310
PV of cash flows PV = CF x DF               1.78               2.38               2.83             51.74
Unlevered value of operations Sum of all PV             58.74

-----------------------------

What is the value of tax shield of SGP?

Year N 1 2 3 4
Tax shield TS               1.00               1.00               2.00               3.00
Horizon value of tax shield HV             32.00
Total Tax shield ITS = TS + HV               1.00               1.00               2.00             35.00
Discount factor (Unlevered cost of equity) Keu 12.20%
Discount factor DF = (1 + Keu)^(-N)           0.8913           0.7944           0.7080           0.6310
PV of cash flows PV = ITS x DF               0.89               0.79               1.42             22.08
Value of tax shield of SGP Sum of all PV             25.19